OpenStudy (anonymous):

What is the slope of a line that passes through the point (−2, 3) and is parallel to a line that passes through (3, 7) and (−2, −8)?

OpenStudy (ajprincess):

First find the slope of the line that passes through (3, 7) and (−2, −8) using \(\large m=\LARGE\frac{y_2-y_1}{x_2-x_1}\)

OpenStudy (anonymous):

can u help me put it together

OpenStudy (anonymous):

Parallel lines have the same slope, so just find the slope of the given points to get your answer. \[m = \frac{ y2 - y1 }{ x2 - x1} \to \frac{ -8 - 7 }{ -2 - 3}\]